A triangular prism is 4cm long, its base is 6cm wide and the volume is 2000cm cubed. what is the height of the prisms base?
Answer by Alan3354(69443)

A triangular prism is 4cm long, its base is 6cm wide and the volume is 2000cm cubed. what is the height of the prisms base?
----------------
V = base area*length
base area = bh/2
V = (bh/2)*l
2000 = (6*h/2)*4
2000 = 12h
h = 2000/12 = 500/3
h =~166.67 cm